OpenStudy (anonymous):

How do i solve mathematical induction dealing with exponents? Such as: 1+7^2+7^3...7^n-1=[(7^n)-1]/6

OpenStudy (anonymous):

do you mean 1 + 7^1 + 7 ^2 ....7^n-1 ?

OpenStudy (anonymous):

yes thats equal to [(7^n)-1]/6

OpenStudy (anonymous):

prove for n=1 \[1= ((7^1)-1)/6\] Assume true for n=k \[1+7 + 7^2 + 7^3...+ 7^k-1 = ((7^k)-1)/6\] prove for n=k+1 \[1+7 + 7^2...+7^k-1 +7^k = (((7^k)-1)/6) + 7^k\] \[= [(7^k)-1)+6(7^k)]/6\] \[=[7(7^k)-1]/6\] =[(7^(k+1))-1]/6

OpenStudy (anonymous):

true for n=1, n=k, n=k+1 therefore true for all values of n where n is a positive integer